# How Do Observable Users Decompose D3 Code? A Qualitative Study

### v4

@@ Abstract
Many toolkit developers seek to streamline the visualization programming process for their users through structured support such as prescribed templates and example galleries. However, few projects examine how users organize their own visualization programs, how their coding choices may deviate from the intents of toolkit developers, and how these differences may impact visualization prototyping and design. Further, is it possible to infer users’ reasoning indirectly through their code, even when users copy code from other sources? Understanding these patterns can reveal opportunities to align toolkit design with actual user behavior, improving usability and supporting more flexible workflows. We explore this question through a qualitative analysis of 715 D3 programs on Observable. We identify three levels of program organization based on how users decompose their code into smaller blocks: Program-, Chart-, and Component-Level code decomposition, with a strong preference for Component-Level reasoning. In a series of interviews, we corroborate that these levels reflect how Observable users reason about visualization programs. We compare common user-made components with those theorized in the Grammar of Graphics to assess overlap in user and toolkit developer reasoning. We find that, while the Grammar of Graphics covers basic visualizations well, it falls short in describing complex visualization types, especially those with animation, interaction, and parameterization components. Our findings highlight how user practices differ from formal grammars and suggest opportunities for rethinking visualization toolkit support, including augmenting learning tools and AI assistants to better reflect real-world coding strategies.
@@ 1 Introduction
In response, we analyze how D3 users organize visualization code to understand what code structures these users find intuitive. Specifically, we apply qualitative methods to investigate an established measurement of code comprehension from the CS education literature: code decomposition [20, 8, 25]. Examining how users decompose code into logical blocks allows us to test implicit assumptions in the literature regarding users’ mental models of visualization programs [18, 4]. Through this analysis, we can establish frameworks for users’ mental models to ground visualization programming and design tools in actual user behavior. Towards this goal, we address two core research questions: How do D3 users (re)organize code copied from outside sources (RQ1)? Further, do D3 users really organize their code according to toolkit designers’ recommendations (RQ2)?
@@
To understand how users’ mental models align with existing toolkit design paradigms (answering RQ2), we compare common user-made D3 components with the Grammar of Graphics (GoG). We find high component overlap for common chart types such as bar and line charts but diminishing coverage for more complex visualizations such as treemaps and streamgraphs. Further, several key D3 component types, such as interaction and parameterization, fall outside the scope of the GoG, suggesting gaps between theory and real-world usage. These insights provide opportunities for building D3 support tools and educational resources that better reflect users’ code organization strategies. In summary, this paper makes the following contributions:
@@ 2 Related Work
Visualization Code Reuse: Copying from examples is a key user strategy for creating D3 visualizations  [6, 2, 4]. For example, in their analysis of 37,815 D3-related posts on Stack Overflow, Battle et al. observe that 14% of them reference just three sources: Observable, the D3 gallery, or Bl.ocks.org  [6]. However, D3 code often contains uncommon syntax and code structures, making it difficult to reuse [6]. In-depth examples may over-complicate prototyping [6] or even lead to design fixation [17]. We seek to understand how users structure their D3 code, which can further clarify how visualization paradigms support user programming in practice. Example galleries are another dominant support for reuse. Yang et al. find that while users want larger galleries, creators struggle to maintain them [30]. Our work complements these findings by studying how example reuse impacts code organization, revealing the impact of galleries on coding structure. Code reuse also has pedagogical value. Recent work showed students who adapted D3 examples created more bespoke visualizations and improved their understanding of the code [12]. This work highlights inherent challenges of adapting D3 code, reinforcing the need to understand how users reason about code in order to better support them.
Visualization Templates: Code reuse is also common in visualization toolkits, which several projects aim to support through code templates. For example, Bako et al. contribute templates for common D3 visualization and interaction types [1]. Harper et al. propose techniques for converting existing D3 visualizations into templates [11]. Tools such as Ivy generalize these ideas to aid creation and reuse [15]. However, templates are difficult to modify beyond their defined parameters, potentially impeding users’ creativity and workflows [4]. Recent work has introduced structured methods for evaluating toolkit notations, reflecting a growing need to systematize how users and developers articulate data transformation and visual mapping [14]. Similarly, our research aims to draw more direct comparisons between users’ and toolkit designers’ organization of visualization programs.
Code Decomposition and Visualization Grammars: Code decomposition is a useful measure of comprehension in CS education [25, 20]. Charitsis et al. find that students who decompose code into modular functions have better assignment outcomes [8]. Decomposition has also been studied in data science notebooks [21]. Raghunandan et al. observe that Jupyter-based notebooks are often decomposed by functionality [19]. Titov et al. even restructure notebooks to improve clarity [26]. These findings align with our study, where users shift decomposition strategies to support iterative thinking. While formalisms such as the Grammar of Graphics [29] and Layered Grammar of Graphics [27] guide toolkit design, it is unclear whether they align with how users reason about code. One study found that users often reuse examples based on layout or task similarity, factors not captured by formal grammars [3]. Analyzing code structure offers a way to bridge this gap [18]. We take a similar approach, focusing on D3 where users still face challenges in writing and debugging visualization programs [6].
@@ 3 Data Collection & Preparation
To facilitate a rigorous qualitative analysis on how users organize their D3 code, we collected a diverse range of 240 Observable [16] notebooks spanning 24 visualization types identified in previous work [5]. We followed a strategy of searching by visualization type and screening notebooks for quality and uniqueness, detailed in supplemental materials (Supplemental Materials).
@@
Further, we define the following terms for our analysis: decompose refers to how users separate code within a single Observable notebook using cells. We use this definition as an overarching term to describe how users “organize,” “structure,” and ”break down” code [9]. Modularity refers to the extent to which users decompose their code into separate pieces, i.e., modules [6, 20]. An example of a module would be an Observable code cell. Sources are D3 programs that provide code for other programs. It could be a notebook on Observable (e.g. from the D3 Gallery) or a D3 program shared externally (e.g. from GitHub Gist).
@@ 4 Can We Infer User Reasoning From D3 Code?
To develop our codebook, a random sample of 15 notebooks from the six most popular D3 visualization types reported in the literature [1, 5, 6] were collected with the procedure in section 3. These 15 notebooks were distinct from the 715 notebooks in the main corpus. The lead authors independently annotated these notebooks, after which the entire author team met to discuss the codebook. The lead authors re-annotated the 15 notebooks again, achieving a Cohen’s Kappa inter-rater reliability score [10] of 0.941. Finally, the lead authors coded the 240 unique notebooks collected from section 3 over 15 weeks. Discrepancies were resolved through discussions. In this section, we detail findings from our coding (the full codebook can be found in our supplemental materials).
Figure 1: Centered are three abstracted notebooks with four color-coded visualization layers. Program-Level has all four layers in one code cell. Chart-Level has a function code cell, a function building code cell, and a data code cell. Lastly, Component-Level has all four layers in different code cells. To the left is a sample of Chart-Level Decomposition. To the right is an example of Component-Level Decomposition.
@@ 4.2 How do D3 Users Leverage Copied Code?
In our codebook, we identify four code-copying strategies; two involve no direct code reuse: “original creation,” where users built notebooks from scratch (56.7%), and “orthogonal code forking,” where users forked a notebook but did not reuse any of the original code (2.1%). The remaining two strategies involve code reuse: “Observable sourced,” where code was reused from another Observable notebook (35.8%), and “outside sourced,” where code was copied from external platforms such as GitHub (5.4%). Over half of the notebooks did not copy code, and among those that did, most sourced from Observable. We analyze how code copying influenced selected decomposition strategies in subsection 4.3.
@@ 4.3 How Does Code Copying Influence Decomposition?
To explore H1 further, we refine our hypothesis: if Observable notebooks are a faithful reflection of the author’s reasoning process, we posit that notebook authors will (1) write code from scratch that reflects their preferred decomposition method or (2) make substantive changes to copied code if the source clashes with how they prefer to decompose D3 code. To test this, we compared decomposition strategies in notebooks with copied code to their sources (i.e., the notebook from which the code was copied). Two key patterns emerged. First, 76.2% of users retained Component-Level decomposition when it was already used in the source. Second, when changing strategies, users shifted exclusively toward Component-Level decomposition. For example, 19.8% shifted from Chart-Level to Component-Level and 9.3% from Program-Level to Component-Level. Further, zero notebooks shifted in the opposite direction.
Thirteen Observable notebooks sourced code from outside platforms that lack Observable’s cell-based environment. Nonetheless, a manual review of their code suggests that 84.62% of the 13 Observable notebooks still employed Component-Level decomposition. These findings suggest that Component-Level decomposition is not simply a byproduct of Observable’s cellular environment and may reflect how users generally reason about D3 code. Similarly, all 7.9% of notebooks with inaccessible deprecated D3 Gallery notebook sources displayed Component-Level decomposition. Lastly, 80.9% of the Original and Orthogonal notebooks used Component-Level decomposition. Together, these results suggest that users prefer Component-Level decomposition, even if they do not inherit any code.
@@ 5.2 Emerging Themes from the Interviews
Using certain decomposition levels also helped participants achieve specific goals. P2 mentioned his team found that using a component approach to organizing the code, instead of using a single cell, led to better visualization rendering in their web application. P3, who used Component-Level decomposition, explained that building up a visualization step-by-step helped him understand the code better. P1 used new cells to enable interactions with his visualizations.
Debugging was a common challenge among participants, in part due to JavaScript’s silent failures [6]. P2 and P3 pointed to splitting code into distinct cells (i.e., Component-Level decomposition) as helping identify problems. P1 and P7 typically do not use Component-Level decomposition, but created new cells to debug. This aligns with CS education work, which finds that Component-Level decomposition can make assignments easier to debug and faster to complete [8].
@@ 6 What Can We Learn From Users’ Decomposition Strategies in D3 Code?
Inspired by prior work [18], we reuse our results from section 4 to compare user-made D3 components with the Grammar of Graphics (GoG) [28] and Layered GoG [27], which formalize how toolkits and languages should ideally be structured. While established formalisms often reflect how toolkit developers think, they do not necessarily reflect how end users themselves reason about visualization code [18]. Thus, we seek to identify and characterize misalignments, which could reveal opportunities to improve toolkit usage and design. Note that we merge overlapping components across the GoG and Layered GoG in our analysis and refer to them collectively as the GoG (Data and Aesthetic Mappings, Statistical Transformation, Geometric Object, Scale, and Coordinate System overlap with the LGoG; the GoG additionally includes Data Transformation).
@@ 6.1 Coverage By Visualization Type
We calculate GoG overlap based on component counts for each visualization type in Table 1. Common visualization types have greater GoG overlap. For example, Battle et al. report Geographic Map, Line Chart, Bar Chart, and Scatterplot as the most popular visualization types in their analysis of D3 usage on the web [5], which we observe also have high overlap with the GoG. However, as complexity increases, GoG overlap drops. This suggests that the Grammar of Graphics supports common charts but is less effective for representing customized designs, where D3 is considered more useful [11, 30], answering RQ2. The diminishing returns between the GoG and D3 suggest an exciting opportunity to infer higher-level abstractions from real-world toolkit usage; in essence, a user-driven generalization of existing template-based approaches. Also, we emphasize that these results generalize to the full dataset of 715 notebooks, including duplicates.
@@ 6.2 Isolated Components
We additionally examined the components that were more commonly isolated in their own cell in Table 2, indicating deliberate structuring decisions (see subsection 4.1). Almost 73% of interactions are isolated, likely due to their complexity, which motivates their separation from other code components [4, 24]. On the other hand, only 54% of animations are separated, a side effect of how D3 is structured. Since animations are often written as calls to existing components, they are harder to reason about and debug. These findings highlight opportunities to support D3 users with new grammars or toolkits that cover GoG components as well as overlooked areas like parameterization, animation, and interaction.
Table 2: Percentage of isolated components across all instances of the component.
@@ 7 Discussion: Implications for Future Research
Analyzing D3 Usage to Assess Existing Theory: To the best of our knowledge, our paper provides the first qualitative analysis of overlap between D3 notebooks and the GoG [28]. For RQ2, we find high alignment for simpler charts like bars and lines, while complex types such as word clouds and Sankey diagrams fall outside the GoG framework. Notably, essential D3 features like animations, interactions, layouts, and parameterization are not covered by the GoG [24]. While the GoG is not expected to capture every use case, our findings offer valuable feedback for rethinking grammar-based language design to better support user needs.
Leveraging Decomposition Strategies for Learning: Our user study shows users purposefully adopt decomposition strategies to support their workflows both when coding from scratch and when reusing code, thus answering RQ1. Educators can leverage our findings by modularizing tutorials or assigning Program-Level code and observing whether students restructure it as an informal assessment. AI tools could also be trained to organize and label code by components (e.g., data handling, interactions) and provide explanations, producing more intuitive, well-structured output.
Developing Resources to Enhance Visualization Design Through Component Reuse: Component-Level decomposition can support efficient design by mapping common visualization components across multiple notebooks and visualization types. A resource linking reusable components across visualizations could help users quickly construct custom charts without starting from scratch. In this way, users can explore D3 code along two levels of abstraction simultaneously: Component-Level semantics and the low-level D3 syntax. Such analysis may even enable automatic component detection in other languages, helping to synthesize language abstraction levels automatically (e.g., generate Vega-Lite analogs for more complex toolkits). By providing a more structured, component-based pathway to visualization design, this method could facilitate faster design iteration and experimentation.
@@
All supplemental materials are available on OSF at https://osf.io/sudb8/?view_only=302fc5c8d397412aac35c6e094ae7dd6. We also provide a complete copy of supplemental materials as a zip file on PCS, with a detailed README.md file that outlines its contents.
@@ Acknowledgements.
The authors wish to thank A, B, and C. This work was supported in part by a grant from XYZ.

### v5

@@ Abstract
Many toolkit developers seek to streamline the visualization programming process through structured support such as prescribed templates and example galleries. However, few projects examine how users organize their own visualization programs how their coding choices may deviate from the intents of toolkit developers, impacting visualization prototyping and design. Further, is it possible to infer users’ reasoning indirectly through their code, even when users copy code from other sources? We explore this question through a qualitative analysis of 715 D3 programs on Observable. We identify three levels of program organization based on how users decompose their code into smaller blocks: Program-, Chart-, and Component-Level code decomposition, with a strong preference for Component-Level reasoning. In a series of interviews, we corroborate that these levels reflect how Observable users reason about visualization programs. We compare common user-made components with those theorized in the Grammar of Graphics to assess overlap in user and toolkit developer reasoning. We find that, while the Grammar of Graphics covers basic visualizations well, it falls short in describing complex visualization types, especially those with animation, interaction, and parameterization components. Our findings highlight how user practices differ from formal grammars and visualization toolkit support, including augmenting learning tools and AI assistants to better reflect real-world coding strategies.
@@ 1 Introduction
In response, we analyze how D3 users organize visualization code to understand what code structures these users find intuitive. We apply qualitative methods to investigate an established measurement of code comprehension : code decomposition [24, 9, 29]. Through this analysis, we can establish frameworks for users’ mental models to ground visualization programming and design tools in actual user behavior. Towards this goal, we address two core research questions: How do D3 users (re)organize code copied from outside sources (RQ1)? Further, do D3 users really organize their code according to toolkit designers’ recommendations (RQ2)?
@@
To understand how users’ mental models align with existing toolkit design paradigms (answering RQ2), we compare common user-made D3 components with the Grammar of Graphics (GoG). We find high component overlap for common chart types such as bar charts but diminishing coverage for more complex such as streamgraphs. Further, several D3 component types, such as interaction and parameterization, fall outside the GoG, suggesting gaps between theory and real-world usage. These insights opportunities for building D3 support tools and resources that better reflect users’ code organization strategies. In summary, this paper makes contributions:
@@ 2 Related Work
Visualization Code Reuse: Copying from examples is a key user strategy for creating D3 visualizations  [6, 2, 4]. For example, in their analysis of 37,815 D3-related posts on Stack Overflow, Battle et al. observe that 14% of them reference just three sources: Observable, the D3 gallery, or Bl.ocks.org  [6]. However, D3 code often contains uncommon syntax and code structures, making it difficult to reuse [6]. In-depth examples may over-complicate prototyping [6] or even lead to design fixation [20]. Example galleries reuse. Yang et al. find that while users want larger galleries, creators struggle to maintain them [35]. Code reuse also has pedagogical value. Recent work showed students who adapted D3 examples created more bespoke visualizations and improved their understanding of the code [13].
Visualization Templates: Several projects aim to support through code templates. For example, Bako et al. contribute templates for common D3 visualization and interaction types [1]. Harper et al. propose techniques for converting existing D3 visualizations into templates [12]. Tools such as Ivy generalize these ideas to aid creation and reuse [18]. However, templates are difficult to modify beyond their defined parameters, potentially impeding users’ creativity and workflows [4]. Recent work has introduced structured methods for evaluating toolkit notations, reflecting a growing need to systematize how users and developers articulate data transformation and visual mapping [17].
Code Decomposition and Visualization Grammars: Code decomposition is measure of comprehension in CS education [29, 24, 34].
Formalisms such as the Grammar of Graphics [33] and Layered Grammar of Graphics [31] toolkit design ). , it is unclear whether they align with how users reason about code. One study found that users often reuse examples based on layout or task similarity, factors not captured by formal grammars [3]. Analyzing code structure offers a way to bridge this gap [21].
@@ 3 Data Collection & Preparation
Further, we define the following terms for our analysis: decompose refers to how users separate code within a single Observable notebook using cells [10]. Modularity refers to the extent to which users decompose their code into separate pieces, i.e., modules [6, 24]. Sources are D3 programs that provide code for other programs. It could be a notebook on Observable (e.g. from the D3 Gallery) or D3 program (e.g. from GitHub Gist).
@@ 4 Can We Infer User Reasoning From D3 Code?
To develop our codebook, a random sample of 15 notebooks from the six most popular D3 visualization types were collected with the procedure in section 3. These 15 notebooks were distinct from the 715 notebooks in the main corpus. The lead authors independently annotated these notebooks, after which the entire author team met to discuss the codebook. The lead authors re-annotated the 15 notebooks again, achieving a Cohen’s Kappa inter-rater reliability score [11] of 0.941. Finally, the lead authors coded the 240 unique notebooks collected from section 3 over 15 weeks. Discrepancies were resolved through discussions. In this section, we detail findings from our coding (the full codebook can be found in our supplemental materials).
Example layouts of the three decomposition levels. Diagram shows three notebook structures: Program-Level combines all data, geometric positional layers, functions, and scales in one cell; Chart-Level separates function-calling into a cell and then function building, scales, and data into another cell. Component-Level places each component in its own cell. Example decomposition code is shown on the left for Chart-Level and on the right for Component-Level.
Figure 1: Centered are three abstracted notebooks with four color-coded visualization components. Program-Level has all four components in one code cell. Chart-Level has a function code cell, a function building code cell, and a data code cell. Lastly, Component-Level has all four components in different code cells. To the left is a sample of Chart-Level Decomposition. To the right is an example of Component-Level Decomposition.
@@ 4.2 How do D3 Users Leverage Copied Code?
We identify four code-copying strategies; two involve no direct code reuse: “original creation,” where users built notebooks from scratch (, 56.7%), and “orthogonal code forking,” where users forked a notebook but did not reuse any code (, 2.1%). The remaining two strategies involve code reuse: “Observable sourced,” where code was reused from another Observable notebook (, 35.8%), and “outside sourced,” where code was copied from external platforms such as GitHub (, 5.4%). Over half of the notebooks did not copy code; of those that did, most sourced from Observable. We analyze how code copying influenced selected decomposition strategies in subsection 4.3.
@@ 4.3 How Does Code Copying Influence Decomposition?
To explore H1 further, we refine our hypothesis: if Observable notebooks are a faithful reflection of the author’s reasoning process, we posit that notebook authors will (1) write code from scratch that reflects their preferred decomposition method or (2) make substantive changes to copied code if the source clashes with how they prefer to decompose D3 code. To test this, we compared decomposition strategies in notebooks with copied code to their sources (i.e., the notebook from which the code was copied). Two key patterns emerged. First, 76.2% of retained Component-Level decomposition when it was already used in the source. Second, when changing strategies, users shifted exclusively toward Component-Level decomposition. For example, 19.8% shifted from Chart-Level to Component-Level and 9.3% from Program-Level to Component-Level. Further, zero notebooks shifted in the opposite direction.
Observable notebooks sourced code from outside A manual review of their code suggests that 84.62% notebooks still employed Component-Level decomposition, These findings suggest that Component-Level decomposition is not simply a byproduct of Observable’s cellular environment and may reflect how users generally reason about D3 code. Similarly, all 7.9% of notebooks with inaccessible deprecated D3 Gallery notebook sources displayed Component-Level decomposition. Lastly, 80.9% of the Original and Orthogonal notebooks used Component-Level decomposition. Together, these results suggest that users prefer Component-Level decomposition, even if they do not inherit any code.
@@ 5.2 Emerging Themes from the Interviews
Using certain decomposition levels also helped participants achieve specific goals. P2 mentioned his team found that using a component approach to organizing the code, instead of , led to better visualization rendering in their web application. P3, who used Component-Level decomposition, explained that building a visualization step-by-step helped him understand the code better. P1 used new cells to enable interactions with his visualizations.
Debugging was a common challenge, in part due to JavaScript’s silent failures [6]. P2 and P3 pointed to splitting code into distinct cells (i.e., Component-Level decomposition) as helping identify problems. P1 and P7 typically do not use Component-Level decomposition, but created new cells to debug. This aligns with work, which finds that Component-Level decomposition can make assignments easier to debug and faster to complete [9].
@@ 6 What Can We Learn From Users’ Decomposition Strategies in D3 Code?
Inspired by prior work [21], we reuse our results from section 4 to compare user-made D3 components with the Grammar of Graphics (GoG) [32] and Layered GoG [31], which formalize how toolkits and languages should ideally be structured. While established formalisms often reflect how toolkit developers think, they do not necessarily reflect how end users themselves reason about visualization code [21]. Thus, we seek to identify and characterize misalignments, which could reveal opportunities to improve toolkit usage and design. Note that we collectively as the GoG (Data and Aesthetic Mappings, Statistical Transformation, Geometric Object, Scale, and Coordinate System overlap with the LGoG; the GoG additionally includes Data Transformation).
@@ 6.1 Coverage By Visualization Type
We calculate GoG overlap based on component counts for each visualization type in Table 1. Common visualization types have greater GoG overlap. For example, Battle et al. report Geographic Map, Line Chart, Bar Chart, and Scatterplot as the most popular visualization types in their analysis of D3 usage [5], which also have high overlap with the GoG. As complexity increases, GoG overlap drops. This suggests that the Grammar of Graphics supports common charts but is less effective for representing customized designs, where D3 is considered more useful [12, 35], answering RQ2. The diminishing returns between the GoG and D3 suggest an exciting opportunity to infer higher-level abstractions from real-world toolkit usage. Also, we emphasize that these results generalize to the full dataset of 715 notebooks, including duplicates.
@@ 6.2 Isolated Components
We examined components that were isolated in their own cell in Table 2, indicating deliberate structuring decisions. of interactions are isolated, likely due to their complexity, which motivates their separation from other components [4, 28]. Only 54% of animations are separated, a likely side effect of how as calls to existing components, harder to reason about and debug. These findings to overlooked areas like parameterization, animation, and interaction.
Table 2: Percentage of .
@@ 7 Discussion: Implications for Future Research
Analyzing D3 Usage to Assess Existing Theory: To the best of our knowledge, our paper provides the first qualitative analysis of overlap between D3 and the GoG [32]. For RQ2, we find high alignment for simpler charts like bars and lines, while complex types such as word clouds and Sankey diagrams fall outside the GoG framework. Notably, essential D3 features like animations, interactions, layouts, and parameterization not covered by the GoG [28]. While the GoG is not expected to capture every use case, our findings
Leveraging Decomposition Strategies for Learning: Our user study shows users purposefully adopt decomposition strategies to support their workflows both when coding from scratch and when reusing code, thus answering RQ1. Educators can leverage our findings by modularizing tutorials or assigning Program-Level code and observing students restructure it as an informal assessment. AI tools could also be trained to organize and label code by components (e.g., data handling, interactions) and provide explanations, producing more intuitive, well-structured output.
Developing Resources to Enhance Visualization Design Through Component Reuse: Component-Level decomposition can support efficient design by mapping common visualization components across multiple notebooks and visualization types. A resource linking reusable components across visualizations could help users quickly construct custom charts without starting from . In this way, users can explore D3 code along two levels of abstraction simultaneously: Component-Level semantics and the low-level D3 syntax. Such analysis may even enable automatic component detection in other languages, helping to synthesize language abstraction levels automatically (e.g., generate Vega-Lite analogs for more complex toolkits). By providing a more structured, component-based pathway to visualization design, this method could facilitate faster design iteration and experimentation.
@@
All supplemental materials are available on OSF at .